# Everybody Prune Now: Structured Pruning of LLMs with only Forward Passes

### v2

@@ 1 Introduction
Table 1: Landscape of resource consumption (memory and compute) of different model compression methods at training time and the inference time resource consumption of the models they deliver. ✗  means the method incurs a prohibitive cost to the lay practitioner whilst ✓  denotes that it is a viable option with respect to that resource.
@@
Quantization (Mixed Precision)
Distillation
Unstructured Pruning
Gradient-Based Structured Pruning
Bonsai (Ours)
@@
Unfortunately, so far, these methods have fallen short of the goal of truly democratizing access to LLMs. In the process of producing a smaller model out of an LLM, methods such as distillation and structured pruning are inaccessible to the everyday practitioner due to their prohibitive resource consumption at training time. Specifically, pure distillation-based techniques require running LLMs to generate large amounts of teacher data (Jiao et al., 2019; Hsieh et al., 2023) whilst existing structured pruning approaches like LLM-Pruner (Ma et al., 2023) and LoRAPrune (Zhang et al., 2023) require several times more memory than is needed to run inference on the model being pruned. Though unstructured pruning (Frantar & Alistarh, 2023; Sun et al., 2023) and quantization are less restrictive at training time, the models they produce are not faster except in the presence of specialized hardware for the former (Mishra et al., 2021), whilst the latter can actually slow down inference due to added overhead (Dettmers et al., 2022). This limits the usefulness of these options for practitioners who are concerned with latency-critical applications. Table 1 summarizes the landscape of existing methods and their limitations.
[…]
The practitioner only has enough memory on their hardware to run inference on the model to be pruned. [^1] Specifically we assume a forward pass with a batch size of at least 1. Given the hardware and open source LLM landscape as of January 2024, for an Nvidia A6000 (48GB) GPU, the largest (full-precision) model this would be applicable to is LLaMA-7B
The above setting is evergreen. As state-of-the-art models become more compute intensive over time, the generational gap in hardware available to the lay practitioner versus the most resource endowed institutions is expected to persist or possibly widen.
In light of the proposed setting, we present Bonsai, a forward pass-only structured pruning approach that is capable of delivering fast, compact, and accurate pruned models under the memory limitations that are typical of consumer hardware. To decide which modules (attention head, rows in feedforward projection, etc.) of the LLM to prune, Bonsai estimates module importances perturbatively by generating sub-models and evaluating their performance (running inference). We make this approach tractable by contributing multiple techniques.
First, we treat the problem of inferring each module’s importance from the performance of generated sub-models as an under-determined regression problem. This enables us to estimate the importance of a large number of modules by exploring a manageable number of random sub-models. This is unlike past perturbative approaches, which prohibitively require roughly as many sub-models as there are modules to select from (Ancona et al., 2020), making them intractable for LLMs. Next, instead of instantiating sub-models by dropping modules with equal likelihood (Kang et al., 2023), we use informative priors derived from work on unstructured pruning (Han et al., 2015; Sun et al., 2023). We thus obtain better estimates of module relevance with fewer evaluated sub-models. Finally, unlike past gradient-free approaches that greedily make pruning decisions layer-by-layer (Dekhovich et al., 2021; Nova et al., 2023; Sun et al., 2023), Bonsai takes a holistic view to preserve the accuracy of the pruned model: modules across layers are removed and evaluated together and relevance scores are computed globally to make pruning decisions. To the best of our knowledge, these ingredients taken together make Bonsai the first successful attempt at scaling gradient-free structured pruning to LLMs.
We conduct several experiments that demonstrate Bonsai’s efficacy. Bonsai, which uses only forward passes for pruning, achieves comparable performance to 2:4 semi-structured sparsity with Wanda (Sun et al., 2023) (0.75×0.75$\times$0.75 × speedup after fine-tuning with LoRA(Hu et al., 2021)) whilst also speeding up inference by up to 1.58×1.58$\times$1.58 × compared to the parent model. Even when compared to SoTA gradient-based structured pruning methods like LLM-Pruner and LoRAPrune, Bonsai outperforms these methods—on 4 out of 6 evaluation settings in our experiments. We also use Bonsai to prune the $\sim$3B Phi-2 (Li et al., 2023) model to a $\sim$1.8B model that outperforms the best sub-2B parameter model on the Huggingface Open LLM leaderboard on 4 out of 6 tasks as of the time of submission. Based on these strong results and its usability under real-world memory constraints, we view Bonsai as a significant contribution to unlocking the power of LLMs for a broader spectrum of practitioners facing diverse hardware constraints.
2 Background on Pruning, Problem Definition and Notation Setup
$\begin{split}\mathbf{m}^{\ast}&=\mathrm{argmax}_{\bar{\mathbf{m}}\in\mathcal{F}_{p}}\quad U\left(\mathbf{M}_{|\bar{\mathbf{m}}}\right)\qquad\text{where}\\ \mathcal{F}_{p}&=\bigg{\{}\bar{\mathbf{m}}\subseteq\mathbf{m}~{}\bigg{|}~{}\bigg{(}\sum_{[j:m_{j}\in\bar{\mathbf{m}}]}s_{j}\bigg{)}\leq(1-p)D\bigg{\}}\end{split}$ (1)
$\mathcal{F}_{p}$ consists of all sub-models that meet the sparsity threshold. Note that, in general, not only does $\mathbf{M}_{|\mathbf{m}^{\ast}}$ have a smaller memory footprint than $\mathbf{M}$, it is also faster to run inference on it since it has fewer modules.
Many structured pruning methods attempt to solve Equation 1 by gradient-guided optimization (or search) over the space of sub-models. However, since we are interested in the memory-constrained setting where computing gradients is not feasible, these methods cannot be used. We will thus focus on developing a memory-friendly structured pruning technique, but we will compare to semi-structured pruning methods since they also have minimal memory overhead.
3 Methodology
In this section, we will develop a structured pruning algorithm that relies exclusively on forward passes through the parent model.
3.1 Estimating module relevance with only forward passes
We have motivated the setting of pruning a model that is so large (relative to the amount of memory available to the practitioner), such that we can only run forward passes through it. This means that we have to solve Equation 1 by relying on only evaluations of $U$, as opposed to gradient-based optimization with respect to $U$. A brute force approach would enumerate all members of $\mathcal{F}_{p}$, find their corresponding performances and pick the best one. Unfortunately, this would be computationally infeasible since $|\mathcal{F}_{p}|$ is combinatorial in the size of the model. [^2] The typical fully connected (FC) sub-layer in an LLM has size $>10^{4}$ units. Pruning just 1 of these FCs to 0.5 sparsity in this way requires evaluating over $\approx$ ($10^{4}\choose 10^{3}$ subsets.
We propose a computationally tractable approach where we first perform a small number, $n$, of evaluations, where $n\ll|\mathcal{F}_{p}|$, to gather data for estimating the relevance of each module in $\mathbf{M}$ with respect to the metric $U$. Upon carrying out this estimation, we can greedily choose the member of $\mathcal{F}_{p}$ that has the highest total module relevance. Specifically, let us assume that we have estimated $\mathbf{\beta}=\{\beta_{i}\}_{i\in[N]}$ to be the relevance of each of the $N$ modules. We can generate an approximate solution to Equation 1 as:
$\mathbf{m}^{\ast}\approx\mathbf{m}^{\mathrm{approx}}=\mathrm{argmax}_{\bar{\mathbf{m}}\in\mathcal{F}_{p}}\quad\sum_{j\in\bar{\mathbf{m}}}\beta_{j}$ (2)
We note that Equation 2 is straightforward to solve, as it simply requires sorting $\beta_{j}$s and greedily selecting the top modules until the parameter count constraint is met (this greedy approach may slightly overshoot the sparsity constraint but since $s_{i}\ll(1-p)D~{}\forall~{}i$ for our settings of interest, the difference is not significant).
Estimating $\mathbf{\beta}$:
To obtain estimates of the module relevance scores $\mathbf{\beta}=\{\beta_{i}\}_{i\in[N]}$, we propose to generate and evaluate $n\ll|\mathcal{F}_{p}|$ sub-models $\bar{\mathbf{m}}_{k}$, and construct a dataset of the sampled sub-models and their corresponding evaluated performances: $\mathbb{D}=\left\{\bar{\mathbf{m}}_{k},U_{k}\right\}_{k\in[n]}$ where $U_{k}=U(\mathbf{M}_{|\bar{\mathbf{m}}_{k}})$. Given $\mathbb{D}$, we treat the problem of estimating $\mathbf{\beta}$ as an under-specified regression problem :
$\begin{split}\hat{\mathbf{\beta}}&=\mathrm{argmin}_{\mathbf{\beta}\in\mathbb{R}^{N}}\quad\frac{1}{n}\sum_{\left(\bar{\mathbf{m}}_{k},U_{k}\right)\in\mathbb{D}}\left(U_{k}-\beta^{T}\alpha_{\bar{\mathbf{m}}_{k}}\right)^{2}+\gamma\|\mathbf{\beta}\|\\ \end{split}$ (3)
where $\alpha_{\bar{\mathbf{m}}_{k}}\in\mathbb{R}^{N}~{}\big{|}~{}\left(\alpha_{\bar{\mathbf{m}}_{k}}\right)_{i}=\mathbf{1}[i\in\bar{\mathbf{m}}_{k}]$, is the binary vector that has 0 at indices where modules have been dropped.
Implementing a sub-model $\bar{\mathbf{m}}_{k}$ as a binary mask $\alpha_{\mathbf{m}_{k}}$ is key to practically realizing our approach. We never actually fully instantiate sub-models as this would be prohibitively expensive. Instead, we create sub-models virtually by zeroing out the outputs of the components to be pruned so they have no effect on the final model output.
3.2 Selecting sub-models for evaluation
An as yet unexplored design choice is how to choose the $n$ candidate sub-models for evaluation. A naive approach would be to sample uniformly from the space of all feasible sub-models. However, we can show that this is sub-optimal. Let us assume $m_{i}$ is a module that is critical for good performance under evaluation with $U$. Since $n<N$, it means that some modules will never be ”turned on” in the list of $n$ chosen sub-models. If $m_{i}$ happens to be one of these masked-out modules, the resulting estimate of $\hat{\beta_{i}}=0$ would in turn result in $\mathbf{m}^{\mathrm{approx}}$ being a poor estimate for $\mathbf{m}^{\ast}$. We therefore need to make a more informed choice to ensure that our estimates of $\mathbf{\beta}$ are accurate and useful.
Given a module $m_{i}$, we can set the likelihood of it being present in any of the $n$ sampled sub-models to be proportional to some measure $\rho_{i}$ which captures our prior belief that $m_{i}$ is a useful module. To define $\rho_{i}$, we can turn to metrics from the unstructured pruning literature. Specifically, we set $\rho_{i}$ to be a module-level analogue of any of the parameter-level unstructured pruning metrics like Wanda (Sun et al., 2023) or output activation magnitude. As an illustrative example, consider that we have shallow network with 1 hidden layer (of dimension $d$) we wish to prune. If output activation magnitude is the metric for our prior, we would calculate the activation vector for the $d$ hidden dimensions – $\hat{\mathbf{a}}$ – averaged across multiple samples. The likelihood that the $i$th column of the matrix $W\in R^{d\times d}$ is left unpruned would be:
$\rho_{i}\propto\hat{\mathbf{a}}_{i}=\frac{1}{B}\sum_{b}\bigg{|}\sigma\bigg{(}\big{(}W^{T}[i,:]\big{)}x_{b}\bigg{)}\bigg{|}$
$\sigma\text{ is the nonlinearity}$. Sampling sub-models according to $\mathbf{\rho}$ means we are more likely to explore models that already have a good performance $U$. The priors $\rho_{i}$ can be computed via forward passes through the original unmodified model $\mathbf{M}_{\theta}$ which allows us to respect memory constraints.
To enhance the efficiency of our method, instead of considering all active modules for pruning, in each layer we consider the bottom $2p$ fraction of modules [^3] The choice of $2p$ is arbitrary. Practitioners can run hyper-parameter searches to find more optimal values for their settings. ranked according to our prior $\rho$. Thus, the top $1-2p$ fraction of modules remain fixed whilst the rest are perturbed and their relevance scores are computed and compared for pruning. This helps reduce the space of possible sub-models being considered for evaluation. For the bottom $2p$ fraction being compared, whenever we generate a mask $\alpha_{\mathbf{m}_{k}}$ with sparsity level $p$, we also generate its complement $\alpha^{c}_{\mathbf{m}_{k}}$, which is obtained by flipping the values in $\alpha_{\mathbf{m}_{k}}$ (except for the fixed $1-2p$ fraction of entries). Covert & Lee (2020) show that this technique can help reduce the variance of the estimator obtained from regression with binary inputs.
@@ 3.2 Selecting sub-models for evaluation
5:     Sample $\{\mathbf{\bar{m}}_{i}\}_{[\frac{n_{\mathrm{iter}}}{2}]}$ sub-models according to $\bar{\mathbb{\rho}}^{l}$
6:     Generate $\{\mathbf{\bar{m}}^{c}_{i}\}_{[\frac{n_{\mathrm{iter}}}{2}]}$ complements of $\{\mathbf{\bar{m}}_{i}\}_{[\frac{n_{\mathrm{iter}}}{2}]}$
@@
7:     Run forward pass on each sub-model and compute $U$. Construct $\mathbb{D}^{l}=\left\{\bar{\mathbf{m}}_{i},U_{i}\right\}_{[\frac{n_{\mathrm{iter}}}{2}]}\bigcup\left\{\bar{\mathbf{m}}^{c}_{k},U_{k}\right\}_{[\frac{n_{\mathrm{iter}}}{2}]}$
8:     $\beta^{\l}\leftarrow\mathrm{Regress}\left(\mathbb{D}^{l}\right)$
@@
9:     $\{m_{\mathrm{pruned}}\}\leftarrow$ sort $\beta^{\l}$ and remove the bottom $k$ modules that correspond to $p_{\mathrm{iter}}$ fraction of the model.

### v3

@@ Appendix B Background: Additional Methods for LLM Compression
Table 7: Landscape of resource consumption (memory and compute) of different model compression methods at training time and the inference time resource consumption of the models they deliver. ✗  $\rightarrow$ a prohibitive cost to the lay practitioner whilst ✓  $\rightarrow$ a viable option with respect to that resource.
@@
Quantization (Mixed Precision) Distillation Unstructured Pruning Gradient-Based Structured Pruning Bonsai (Ours)
@@
Distillation:
Originally proposed by (Hinton et al., 2015; Sanh et al., 2019) have shown that it is possible to reach similar performances on many downstream tasks using much smaller language models pre-trained with knowledge distillation. The resulting models are lighter and faster at inference time. Unfortunately, as noted by (Xia et al., 2022b), distilling a pre-trained model to a student model that has been initialized from scratch requires lots of compute and data to recover reasonable levels of teacher performance. To reduce this resource burden, (Sanh et al., 2020; Lagunas et al., 2021; Xia et al., 2022b) combine distillation with gradient-based structured pruning, so that the student model does not have to be initialized from scratch. Bonsai also leverages distillation, but only during the post-pruning adaptation phase. This maintains the constraint that the pruning process is gradient-free. And since the pruned model is smaller than the parent model, we can safely perform distillation and parameter efficient fine-tuning (Hu et al., 2021; He et al., 2021) within the bounds of available memory.
Quantization:
(Li et al., 2020) introduced a training strategy that showed that heavily compressed, large models achieved higher accuracy than lightly compressed, small models, providing both efficiency and high-accuracy results on NLP tasks at once. (Dettmers et al., 2022) introduced a new method called ‘LLM.int8()‘, which allows loading large models with 16 or 32-bit weights and using them immediately for inference without any performance degradation by combining vector-wise quantization with mixed-precision decomposition. (Xiao et al., 2023) enabled 8-bit weight, 8-bit activation (W8A8) quantization for LLMs with >100B parameters, which smoothed activation outliers by migrating the quantization difficulty from activations to weights with a mathematically equivalent transformation. Quantization tends to be complementary to pruning, allowing independent exploration and subsequent combinations of both techniques. This paper focuses on gradient-free structured pruning. We leave combining Bonsai with quantization for future work.
Appendix C Main Experiment Details
C.1 Full Algorithm
@@
5:     Sample $\{\mathbf{\bar{m}}_{i}\}_{[n_{\mathrm{iter}}]}$ sub-models according to $\bar{\mathbb{\rho}}^{l}$
@@
6:     Run forward pass on each sub-model and compute $U$. Construct $\mathbb{D}^{l}=\left\{\bar{\mathbf{m}}_{i},U_{i}\right\}_{[n_{\mathrm{iter}}]}$
7:     $\beta^{\l}\leftarrow\mathrm{Regress}\left(\mathbb{D}^{l}\right)$
@@
8:     $\{m_{\mathrm{pruned}}\}\leftarrow$ sort $\beta^{\l}$ and drop the bottom $k$ modules that make up $p_{\mathrm{iter}}$ fraction of the model.